Shift Summary - Day August 25, 1999

Douglas A. Jensen (
Wed, 25 Aug 1999 15:43:38 -0500 (CDT)

Shift Summary for Day Shift, 08/25/99

On Shift: D Jensen, J Wang, Jason L covering for DAJ for a couple hours
while he goes off to a meeting. Sydney, TRD guys, others around

Hours of beam available at all intensities: 8:00/8:30
Hours of beam available at nominal intensity: 7:30/8:30
Hours of E832 data (special runs not included): 7:00/8:30
Hours of special runs (laser scans, muon runs...): 0/8:30

Accounting of lost time:
Machine problems ( Linac, MI ), for some hours, missed spills
Tape changes, various KFRIENDS and other resets.

Runs ended this shift:
Run 14322
Run 14323- BEAM_832 (in progress)

Accesses during shift: none

Overview of shift:
E832 datataking (SEM=6.8E12)

Description of Shift:

run spill time comment

14322 6:05 Old crew gone, new crew here
6.6E12, BOA = 1.16 OK
14322 412 8:07 Tapes full, run ends

14323 1 8:30 New run starts, with 2 KFRIENDS

There are occasional 'Mesage from Pipeline'
windows. Sydney says if CsI monitor plots
look ok, ignore 'em. ( no messages )
60 9:50 Missed 5 or 6 spills. Have been loosing about
10% of the spills. MCR says Linic, MI problems

70 10:01 DYC crate 53 full
50-74 Beam breaking up
126 11:17 Pipeline error, init pipeline crate 13, lost 6 spills
148 11:47 beam will be off for power supply swap
162 noon beam back on
200 1 Trigger hung, ran KFRIEND a couple times. OK
216 13:20 CsI too wide - Val Message. ch 181. Went away
after 3-4 spills.
14:40 get message for Sydney - power supply fixed,
Page Sydney, give him message. Sydney happy!
296 15:05 "Pipeline Errors,please reset pipeline"
we do it, then throw in a few KFRENDS too.
loose 3,4 spills. Then re-roll along
15:30 Next crew here. We wish them many happy
Follow up Notes on Previous Shift
Sydney says split issue mentioned prev shift - it is
a timing/sequencing issue among allocation processes. OK

Items to Follow up on Next Shift